# High-Fidelity Data-Driven Dynamics Model for Reinforcement Learning-based Magnetic Control in HL-3 Tokamak

### v1

@@ 2.4.2 RL-based Control for Zero-Shot Adaptation to Changed Triangularity Target
[…]
Fig. 9: Zero-shot triangularity control in shot #6698. a, The upper and lower triangular control results of shot #6698. Compared to the target used for RL training (black dashed line), the target for this discharge (red dashed line) has a triangularity decreased by $0.2$. The red solid line illustrates the experimental result, and the blue dashdot line represents the post-experiment replay of our dynamics model. b, Plasma LCFS reconstructed from six plasma shape and position parameters ($a$, $\kappa$, $\delta_{l}$, $\delta_{u}$, $R$, $Z$) during the RL control phase.
@@
In shot #6698, similar to the procedure described in Section 2.4.1, RL assumes the discharge control task from PID from $1600$ ms to $1700$ ms. Fig. 9a illustrates the corresponding control results of the upper and lower triangularity. Both the lower triangularity $\delta_{l}$ and the upper triangularity $\delta_{u}$ successfully reach the new target values. Specifically, the lower triangularity $\delta_{l}$ swiftly progresses towards the target following a transition period of approximately $10$ ms, finally reaching it around $1650$ ms. Meanwhile, the upper triangularity $\delta_{u}$ follows a gradual trajectory toward the target during the initial $10$-ms transition phase, accelerates its adjustment after around $20$ ms, and aligns with the target at approximately $1650$ ms. It then maintains alignment with the target trajectory for the remaining $50$ ms. These zero-shot triangularity control results verify the versatility of RL towards effective control in previously unencountered scenarios. Furthermore, the post-experiment replay of the dynamics model from $1600$ ms to $1700$ ms reveals a consistent trend, underscoring the model’s generalizability and robustness for supporting RL in exploratory applications.
3 Discussion
@@
Fig. 10: Comparison of the training and deployment between a, MPC with online, limited-term dynamics model autoregression and b, PPO interacting with dynamics model for training only.
Our work demonstrates the feasibility of data-driven, end-to-end learning and control of plasma current and shape. By avoiding reliance on first-principle-based simulators [9] and auxiliary controlling commands from historical data or controllers [12, 11, 5, 36, 37, 35], our approach significantly differs from existing works. High-dimensional control over a learned model falls into the general scope of model-based reinforcement learning (MBRL) [24]. In this regard, as in Fig. 1b, instead of evolving the plasma state, Ref. [5] mainly uses the output of the learned dynamics model to generate a competent reward function, which limits its applicability to localized reactive control only. Meanwhile, as in Fig. 10, Ref. [35] conducts a receding (limited-term) horizon control through computation-intensive MPC, which involves cumbersome online autoregression of the dynamics model for trajectory sampling, making real-time deployment challenging [56]. On the contrary, as in Fig. 1c, our work provides a proof-of-concept study on long-term high-dimensional dynamics model learning, which further enables the training and deployment of the independent PPO actor for complex magnetic control.
We also note that our work is still in the early stages and owing to limited experimental opportunities, a more comprehensive validation is yet to be carried out. Therefore, further experiments and fine-tuning can be explored for more useful applications. For example, more advanced learning and control objectives involving plasma instability and disruption avoidance [57, 58, 59] can be studied. Additionally, progress in AI fields including advanced DNN design with improved capabilities in time-varying dynamics learning [60, 61, 62, 63, 64] and decision-making [65, 66], and orthogonal ingredients from MBRL to cope with out-of-distribution (OOD) issues [67, 68, 69] can be leveraged. Besides, transfer learning across tokamak devices [70] warrants further investigation. This could verify the potential advantages of a fully data-driven method over first-principle-based simulators, which require sophisticated device adaptation. We believe that as a meaningful extension to broaden the capability of data-driven approaches in the fusion area, our approach promises to more efficiently provide qualified solutions to daily discharge operations in future tokamak devices.
@@ 4.1 HL-3
The HL-3 (previously named HL-2M) Tokamak, an experimental fusion device constructed at the Southwestern Institute of Physics in China, aims to advance the comprehension of fundamental physics and engineering principles crucial for the success of global fusion projects like ITER. HL-3 boasts a plasma current capacity ranging from $1.5$ to $3$ MA and operates within a magnetic field strength of $2.3$ T. Its geometric parameters consist of a major radius of $1.78$ m, a minor radius of $0.65$ m, and an aspect ratio of approximately $2.8$, reflecting its toroidal structure. As shown in Fig. 1a, the poloidal field configuration of HL-3 is intricately managed by a Central Solenoid (CS) coil, complemented by an ensemble of eight pairs of poloidal field (PF) coils strategically distributed as inner coils (PF1-PF4), top-bottom coils (PF5-PF6), and outer coils (PF7-PF8), with the latter trio playing a pivotal role in establishing the divertor configuration, hence their nomenclature as divertor coils [71, 72, 73]. In particular, HL-3 directly manipulates the voltage of $17$ coils, comprising of $1$ CS coil and $16$ PF coils. This manipulation indirectly influences the coil currents and regulates the magnetic field within the vacuum chamber, enabling the confinement and discharge of plasma.
@@ 4.2 Data Pre-Processing
Table 1: Flat-top plasma current distribution in the dataset.
Plasma Current (kA) $300$ $400$ $500$ $600$ $700$ $800$ $1000$
Number of Shot $258$ $94$ $245$ $77$ $97$ $32$ $29$
@@
Typically, the plasma discharge process can be divided into three distinct phases (i.e., ramp-up, flat-top, and ramp-down phases). However, for the ramp-up phase’s plasma, the relatively low current and the difficulty in calculating the eddy currents in passive conductor structures (such as the vacuum chamber and the first wall) often result in inaccurate shape calculations, weakening the correlation between the coil voltage commands and plasma shape. Therefore, considering that the primary objective of the dynamics model is to capture the complex evolution of the plasma state and ultimately facilitate efficient RL training, model training is exclusively based on data from the ramp-up phase where the plasma current surpasses $100$ kA and flat-top phase with a duration of at least $500$ ms. Based on the historical discharge logs of HL-3 across campaigns from 2022 to 2024, after such a meticulous selection procedure, a dataset comprising $832$ successful discharges has been curated, with a temporal resolution of $1$ ms, totaling $1,833,516$ time slices. The categorization of these $832$ shots by flat-top plasma current is detailed in Table 1.
Table 2: Input and output variables used in dynamics model, with the length of the time series $n=30$ and the temporal resolution $\Delta{t}=1$ ms. As the CS coil is composed of two parallel connected sets of coils, the dimension of coil current ($I_{c}$) equals that of coil voltage ($U$) plus $1$.
Input variable Unit Number of channels Description Time
$I_{p}$ kA 1 Plasma current ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$a$ cm 1 Minor radius ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$\kappa$ dimensionless 1 Elongation ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$\delta_{u}$ dimensionless 1 Upper triangularity ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$\delta_{l}$ dimensionless 1 Lower triangularity ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$R$ cm 1 Radial position of plasma geometric center ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$Z$ cm 1 Vertical position of plasma geometric center ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$I_{c}$ A 18 Coil current ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$U$ V 17 Coil voltage ($t-(n-1)\Delta{t}$$\sim$($t$)
$TF$ T 1 Toroidal magnetic field ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
$Vloop$ V 1 Loop voltage ($t-n\Delta{t}$$\sim$($t-\Delta{t}$)
Output variable Unit Number of channels Description Time
$I_{p}$ kA 1 Plasma current $t$
$a$ cm 1 Minor radius $t$
$\kappa$ dimensionless 1 Elongation $t$
$\delta_{u}$ dimensionless 1 Upper triangularity $t$
$\delta_{l}$ dimensionless 1 Lower triangularity $t$
$R$ cm 1 Radial position of plasma geometric center $t$
$Z$ cm 1 Vertical position of plasma geometric center $t$
$I_{c}$ A 18 Coil current $t$
@@
Acknowledging the importance of the historical plasma state and coil commands, the dynamics model integrates a time series as its input to better capture the plasma evolution towards equilibrium. Specifically, given the $1$-kHz control frequency, the model uses a $30$-ms historical dataset for single-step prediction of plasma current, shape and position parameters, as well as the coil currents. A summarized list of the input and output variables of the dynamics model can be found in Table 2. Briefly, this model learns up to $25$ variables related to plasma current, shape, and position, as well as coil current from a $44$-dimensional input, which includes autoregressive predictions of these $25$ variables, $17$ executed coil voltage configurations, the toroidal magnetic field, and loop voltage. In addition, $n=30$ represents the length of the time series, and $\Delta{t}=1$ ms is the temporal resolution. Notably, as the CS coil is composed of two parallel connected sets of coils, the dimension of coil current ($I_{c}$) equals that of coil voltage ($U$) plus $1$.
@@
Throughout the discharge experiments in HL-3, engineering conditions change continuously. As a result, we utilize the most recent discharges for testing while earlier datasets are used for training. This methodology contributes to learning from updated data distributions and accommodates to latest discharge environment. To prevent model overfitting, we designate $50$ discharges from Shot #5893 to Shot #6055 as the validation set, and $70$ discharges, including Shots #6056-#6265 and #5095-#5105, as the test set. Note that the range from Shot #5095 to #5105 corresponds to $10$ discharges of advanced divertor configuration, which exhibit more pronounced morphological variations than standard discharges. During training, a $15$-ms smoothing process is implemented on the training data to effectively mitigate non-electromagnetic interference. Since the training dataset suffers from a paucity of similar data instances, the corresponding results in Fig. 3 could partially validate the accuracy and extrapolation capability of the dynamics model.
@@ 4.3.1 DNN Structure for Dynamics Model
[…]
Fig. 11: The testing results of response model with different structures on the same test dataset. TRF: Transformer; All: Scheduled Sampling + Ensemble + Self-Attention + Adaptive Weight.
Due to the practical demands, plasma discharges are often carried out with identical or similar parameter sets to replicate specific physical outcomes, resulting in a lack of diversity in the dataset. Therefore, although the Transformer excels in parallel computation and can apprehend holistic information owing to its self-attention mechanism [46], it faces overfitting issues under limited and narrow data distribution. The testing MAE for both the Transformer and LSTM networks, as shown in Fig. 11, suggests that under the same training methods, dynamics models utilizing LSTM networks typically outperform those based on the Transformer. As a result, we opt for LSTM as the backbone of our dynamics model.
We utilize a stack of DNNs including a three-layer LSTM network, a self-attention module, and a three-layer fully connected MultiLayer Perceptron (MLP). Specifically, each LSTM layer comprises $128$ hidden neurons with a dropout rate of $0.2$. For the $30$ consecutive steps’ $44$-dimensional input, as in Table 2, the third LSTM layer produces an intermediate hidden layer output with dimensions of $30\times 128$. Subsequently, the inner product between each $128$-length tensor and the last one is computed, followed by a softmax-based normalization operation to scale the values between 00 and $1$. These self-attention values serve as attention weights to aggregate the original LSTM outputs as a $128$-length vector. The integration of a self-attention mechanism into the LSTM allows the assignment of distinctive weights to the intermediate output, rather than placing equal emphasis and neglecting the inter-differences. Fig. 3a validates the effectiveness of this modification. Afterward, the self-attention-processed vector undergoes three MLPs with $128$, $64$, and $25$ neurons and a ReLU activation function, culminating in the single-step prediction of $25$ variables as specified in Table 2.
@@ 4.3.2 DNN Structure for EFITNN
EFITNN [38] is a surrogate model trained with input and output provided by an offline Equilibrium FIT (EFIT) code [39, 40]. It accepts input signals from poloidal array of pick-up coils, flux loops, and coil currents, and concatenates them into a $68$-element vector. EFITNN supports the prediction of $8$ scalar values including six plasma shape parameters as well as safe factor in 95% minor radius $q_{95}$ and plasma inductance $l_{i}$. Besides, it can produce predictions for poloidal magnetic flux distribution $\Psi_{\text{rz}}$ and toroidal current density distribution $J_{p}$ with a resolution up to $129\times 129$. For these heterogeneous outputs, EFITNN adopts a shared $4$-layer MLPs. Subsequently, for scalar prediction, another $4$-layer MLPs with residual connections [54] are leveraged, while $\Psi_{\text{rz}}$ and $J_{p}$ relies on deconvolutional networks [74]. Note that each MLP layer consists of $512$ neutrons with a GeLU activation function. EFITNN adopts simultaneous training for this multi-task learning, and we kindly refer to [38] for the training details. Fig. 8 manifests the prediction accuracy after training. In addition, if only plasma shape and position parameters are predicted, EFITNN, without the use of deconvolutional networks, consumes $0.08$ ms on NVIDIA A100 Tensor Core GPU with TensorRT, fully supporting the real-time deployment.
@@ 4.4.1 Scheduled Sampling
[…]
Fig. 12: Schematic illustration of scheduled sampling. a, Without any preventive training measures, the autoregressive nature of the dynamics models implies a gradually larger deviation between the autoregressive output (blue line) and the true data (orange line), due to the compounding error. b, The scheduling sampling technique exposes the model to real data beyond the model output, with a probabilistic teacher-forcing ratio $p$, during the training.
Without any preventive training measures, the autoregressive nature of dynamics models leads to a gradually increasing deviation between the actual data and model predictions due to compounding error, as illustrated in Fig. 12a. Directly applying such a model in practice can result in exposure bias [75], incurring reduced efficacy and severe misalignment between the training and testing modalities. Therefore, during our training, we incorporate scheduled sampling [76, 77] to stochastically determine the input, by selecting between the model’s past predictive outputs and the current-step actual data according to a teacher-forcing ratio $p\in[0,1]$. In other words, as the ratio $p$ approaches $1$, a greater proportion of the real input data will be used. Conversely, as $p$ turns to 00, the model transitions to a strictly autoregressive mode, solely relying on the preceding steps’ predictions. During the training, we initialize the ratio $p$ to $1$ and linearly decay it to 00 over $300$ epochs. This approach ensures rapid convergence during training and facilitates a seamless shift towards an autoregressive scheme for training RL.
[…]
Fig. 13: The advantage of scheduled sampling over a noise layer. SS: Scheduled Sampling; LSTM w/ $(0.1\sigma$/$0.2\sigma$/$0.3\sigma)$ represents the combination of LSTM and a zero-mean Gaussian noise layer with a standard deviation of $0.1\sigma$/$0.2\sigma$/$0.3\sigma$.
In contrast, [11] adopts a different approach to intentionally add a zero-mean Gaussian noise layer to the training data. Nevertheless, apart from the conflict with the data smoothing operation during training, the comparison in Fig. 13 also suggests its inferiority to scheduled sampling. Specifically, the diverse impact on high-dimensional multi-task learning makes selecting the appropriate variance of added Gaussian noise challenging. For example, with an amplitude of $0.1\sigma_{x}$, where $\sigma_{x}$ indicates the standard deviation of the input data, the model exhibits the lowest error in predicting $\kappa$ but the highest error in predicting $I_{p}$. Conversely, the incorporation of scheduled sampling maintains stability in significantly reducing prediction errors for $25$ variables. Therefore, we primarily take account of scheduled sampling in this paper.

### v2

@@ 2.4.2 RL-based Control for Zero-Shot Adaptation to Changed Triangularity Target
Fig. 9: Zero-shot triangularity control in shot #6698. The upper and lower triangular control results of shot #6698. Compared to the target used for RL training (black dashed line), the target for this discharge (green dashed line) has a triangularity decreased by $0.2$. The red solid line illustrates the experimental result, and the blue dashdot line represents the post-experiment replay of our dynamics model.
Table 2: Triangularity Parameters during RL Control for Shot #6698
Time (ms) Upper Triangularity ($\delta_{\mathrm{u}}$) Lower Triangularity ($\delta_{\mathrm{l}}$)
1610 0.235 0.589
1630 0.224 0.597
1640 0.220 0.577
1650 0.189 0.482
1660 0.144 0.412
1690 0.131 0.423
Target 0.128 0.396
@@
In shot #6698, similar to the procedure described in the preceding subsection on target tracking control, RL assumes the discharge control task from PID from $1600$ ms to $1700$ ms with $10$ ms transition periods at the start and end. Fig. 9 and Table 2 illustrate the corresponding control results of the upper and lower triangularity. Both the lower triangularity $\delta_{\mathrm{l}}$ and the upper triangularity $\delta_{\mathrm{u}}$ successfully reach the new target values. Specifically, the lower triangularity $\delta_{\mathrm{l}}$ swiftly progresses towards the target following a transition period of approximately $10$ ms, finally reaching it around $1650$ ms. Meanwhile, the upper triangularity $\delta_{\mathrm{u}}$ follows a gradual trajectory toward the target during the initial $10$-ms transition phase, accelerates its adjustment after around $20$ ms, and aligns with the target at approximately $1650$ ms. It then maintains alignment with the target trajectory for the remaining $50$ ms. These zero-shot triangularity control results verify the versatility of RL towards effective control in previously unencountered scenarios. Furthermore, the post-experiment replay of the dynamics model from $1600$ ms to $1700$ ms reveals a consistent trend, underscoring the model’s generalizability and robustness for supporting RL in exploratory applications.
3 Conclusions
@@
Fig. 10: Comparison of the training and deployment between MPC and PPO. a, MPC with online, limited-term dynamics model autoregression and b, PPO interacting with dynamics model for training only.
Our work demonstrates the feasibility of data-driven, end-to-end learning and control of plasma current and shape. By avoiding reliance on first-principle-based simulators [9] and auxiliary controlling commands from historical data or controllers [12, 11, 5, 36, 37, 35], our approach significantly differs from existing works. High-dimensional control over a learned model falls into the general scope of model-based reinforcement learning (MBRL) [24]. In this regard, as in Supplementary Figure 1a, instead of evolving the plasma state, Ref. [5] mainly uses the output of the learned dynamics model to generate a competent reward function, which limits its applicability to localized reactive control only. Meanwhile, as in Fig. 10, Ref. [35] conducts a receding (limited-term) horizon control through computation-intensive MPC, which involves cumbersome online autoregression of the dynamics model for trajectory sampling, making real-time deployment challenging [56]. On the contrary, as in Supplementary Figure 1b, our work provides a proof-of-concept study on long-term high-dimensional dynamics model learning, which further enables the training and deployment of the independent PPO actor for complex magnetic control.
Our research presents a promising direction for routine discharge operations in future tokamak devices. In the future, we will focus on further improving the overall performance and generalization capability of the control model. Building upon this foundation, we aim to address more challenging control tasks, including the intelligent avoidance of plasma instabilities and disruptions [57, 58, 59]. Moreover, we plan to conduct transfer learning across tokamak devices [60], further verifying the potential advantages of a fully data-driven method over first-principle-based simulators.
@@ 4.1 HL-3
The HL-3 (previously named HL-2M) Tokamak is an experimental fusion device constructed at the Southwestern Institute of Physics in China. It is designed to operate at plasma current $I_{\mathrm{p}}$ =33 MA, toroidal field $B_{\mathrm{t}}=3$ T, major radius $R=1.78$ m, minor radius $a=0.65$ m, elongation $\kappa\leq 1.8$, triangularity $\delta>0.5$ [61]. HL-3 has recently focused on high-current, high $\beta_{\mathrm{N}}$ discharges to achieve high performance plasma, establishing critical foundations for burning plasma research. As shown in Fig. 1a, the poloidal field configuration of HL-3 is intricately managed by a Central Solenoid (CS) coil, complemented by an ensemble of eight pairs of Poloidal Field (PF) coils strategically distributed as inner coils (PF1-PF4), top-bottom coils (PF5-PF6), and outer coils (PF7-PF8), with the latter trio playing a pivotal role in establishing the divertor configuration, hence their nomenclature as divertor coils [62, 63, 64]. In particular, HL-3 directly manipulates the voltage of $17$ coils, comprising of $1$ CS coil and $16$ PF coils. This manipulation indirectly influences the coil currents and regulates the magnetic field within the vacuum chamber, enabling the confinement and discharge of plasma.
@@ 4.2 Data Pre-Processing
Typically, the plasma discharge process can be divided into three distinct phases (i.e., ramp-up, flat-top, and ramp-down phases). However, for the ramp-up phase’s plasma, the relatively low current and the difficulty in calculating the eddy currents in passive conductor structures (such as the vacuum chamber and the first wall) often result in inaccurate shape calculations, weakening the correlation between the coil voltage commands and plasma shape. Therefore, considering that the primary objective of the dynamics model is to capture the complex evolution of the plasma state and ultimately facilitate efficient RL training, model training is exclusively based on data from the ramp-up phase where the plasma current surpasses $100$ kA and flat-top phase with a duration of at least $500$ ms. Based on the historical discharge logs of HL-3 across campaigns from 2022 to 2024, after such a meticulous selection procedure, a dataset comprising $832$ successful discharges has been curated, with a temporal resolution of $1$ ms, totaling $1,833,516$ time slices. The categorization of these $832$ shots by flat-top plasma current is detailed in Supplementary Table 1.
@@
Acknowledging the importance of the historical plasma state and coil commands, the dynamics model integrates a time series as its input to better capture the plasma evolution towards equilibrium. Specifically, given the $1$-kHz control frequency, the model uses a $30$-ms historical dataset for single-step prediction of plasma current, shape and position parameters, as well as the coil currents. A summarized list of the input and output variables of the dynamics model can be found in Supplementary Table 2. Briefly, this model learns up to $25$ variables related to plasma current, shape, and position, as well as coil current from a $44$-dimensional input, which includes autoregressive predictions of these $25$ variables, $17$ executed coil voltage configurations, the toroidal magnetic field, and loop voltage. In addition, $n=30$ represents the length of the time series, and $\Delta{t}=1$ ms is the temporal resolution. As the CS coil is composed of two parallel connected sets of coils, the dimension of coil current ($I_{\mathrm{c}}$) equals that of coil voltage ($U$) plus $1$.
@@
Throughout the discharge experiments in HL-3, engineering conditions change continuously. As a result, we utilize the most recent discharges for testing while earlier datasets are used for training. This methodology contributes to learning from updated data distributions and accommodates to latest discharge environment. To prevent model overfitting, we designate $50$ discharges from shot #5893 to shot #6055 as the validation set, and $70$ discharges, including shots #6056-#6265 and #5095-#5105, as the test set. Note that the range from shot #5095 to #5105 corresponds to $10$ discharges of advanced divertor configuration, which exhibit more pronounced morphological variations than standard discharges. During training, a $15$-ms smoothing process is implemented on the training data to effectively mitigate non-electromagnetic interference. Since the training dataset suffers from a paucity of similar data instances, the corresponding results in Supplementary Figure 4 could partially validate the accuracy and extrapolation capability of the dynamics model.
@@ 4.3.1 DNN Structure for Dynamics Model
Due to the practical demands, plasma discharges are often carried out with identical or similar parameter sets to replicate specific physical outcomes, resulting in a lack of diversity in the dataset. Therefore, although the Transformer excels in parallel computation and can apprehend holistic information owing to its self-attention mechanism [46], it faces overfitting issues under limited and narrow data distribution. The testing MAE for both the Transformer and LSTM networks, as shown in Supplementary Figure 2, suggests that under the same training methods, dynamics models utilizing LSTM networks typically outperform those based on the Transformer. As a result, we opt for LSTM as the backbone of our dynamics model.
We utilize a stack of DNNs including a three-layer LSTM network, a self-attention module, and a three-layer fully connected MultiLayer Perceptron (MLP). Specifically, each LSTM layer comprises $128$ hidden neurons with a dropout rate of $0.2$. For the $30$ consecutive steps’ $44$-dimensional input, as in Supplementary Table 2, the third LSTM layer produces an intermediate hidden layer output with dimensions of $30\times 128$. Subsequently, the inner product between each $128$-length tensor and the last one is computed, followed by a softmax-based normalization operation to scale the values between 0 and $1$. These self-attention values serve as attention weights to aggregate the original LSTM outputs as a $128$-length vector. The integration of a self-attention mechanism into the LSTM allows the assignment of distinctive weights to the intermediate output, rather than placing equal emphasis and neglecting the inter-differences. Fig. 3a validates the effectiveness of this modification. Afterward, the self-attention-processed vector undergoes three MLPs with $128$, $64$, and $25$ neurons and a ReLU activation function, culminating in the single-step prediction of $25$ variables as specified in Supplementary Table 2.
@@ 4.3.2 DNN Structure for EFITNN
EFITNN [38] is a surrogate model trained with input and output provided by an offline Equilibrium FIT (EFIT) code [39, 40]. It accepts input signals from poloidal array of pick-up coils, flux loops, and coil currents, and concatenates them into a $68$-element vector. EFITNN supports the prediction of $8$ scalar values including six plasma shape parameters as well as safe factor in 95% minor radius $q_{\mathrm{95}}$ and plasma inductance $l_{\mathrm{i}}$. Besides, it can produce predictions for poloidal magnetic flux distribution $\Psi_{\mathrm{rz}}$ and toroidal current density distribution $J_{\mathrm{p}}$ with a resolution up to $129\times 129$. For these heterogeneous outputs, EFITNN adopts a shared $4$-layer MLPs. Subsequently, for scalar prediction, another $4$-layer MLPs with residual connections [54] are leveraged, while $\Psi_{\text{rz}}$ and $J_{p}$ relies on deconvolutional networks [65]. Note that each MLP layer consists of $512$ neutrons with a GeLU activation function. EFITNN adopts simultaneous training for this multi-task learning, and we kindly refer to [38] for the training details. Fig. 8 manifests the prediction accuracy after training. In addition, if only plasma shape and position parameters are predicted, EFITNN, without the use of deconvolutional networks, consumes $0.08$ ms on NVIDIA A100 Tensor Core GPU with TensorRT, fully supporting the real-time deployment.
@@ 4.4.1 Scheduled Sampling
Without any preventive training measures, the autoregressive nature of dynamics models leads to a gradually increasing deviation between the actual data and model predictions due to compounding error, as illustrated in Supplementary Figure 3. Directly applying such a model in practice can result in exposure bias [66], incurring reduced efficacy and severe misalignment between the training and testing modalities. Therefore, during our training, we incorporate scheduled sampling [67, 68] to stochastically determine the input, by selecting between the model’s past predictive outputs and the current-step actual data according to a teacher-forcing ratio $p\in[0,1]$. In other words, as the ratio $p$ approaches $1$, a greater proportion of the real input data will be used. Conversely, as $p$ turns to 0, the model transitions to a strictly autoregressive mode, solely relying on the preceding steps’ predictions. During the training, we initialize the ratio $p$ to $1$ and linearly decay it to 0 over $300$ epochs. This approach ensures rapid convergence during training and facilitates a seamless shift towards an autoregressive scheme for training RL.
In contrast, [11] adopts a different approach to intentionally add a zero-mean Gaussian noise layer to the training data. Nevertheless, apart from the conflict with the data smoothing operation during training, the comparison in Supplementary Figure 4 also suggests its inferiority to scheduled sampling. Specifically, the diverse impact on high-dimensional multi-task learning makes selecting the appropriate variance of added Gaussian noise challenging. For example, with an amplitude of $0.1\sigma_{x}$, where $\sigma_{x}$ indicates the standard deviation of the input data, the model exhibits the lowest error in predicting $\kappa$ but the highest error in predicting $I_{\mathrm{p}}$. Conversely, the incorporation of scheduled sampling maintains stability in significantly reducing prediction errors for $25$ variables. Therefore, we primarily take account of scheduled sampling in this paper.